A 106 g Frisbee is 27 cm in diameter and has half
Answer #1

I = Iring + Idisk

= (m/2)*r^2 +  (m/2)*r^2 / 2

= ( 0.106/2) *0.27^2 + ( 0.106/2) *0.27^2 /2

I = 0.005795 kg .m^2

w = 510 rpm = 510 * 2pi / 60 rad/s = 53. 41 rad/s

a(t) = 1/2* \alpha t^2
and the angular velocity
w = \alpha * t ........................1

pi/2 = 1/2 \alpha ​t^2

pi =  \alpha*t^2 ................2

from 1

53. 41 = \alpha*t

\alpha = 53.41 /t

put in 2

pi = (53.41 /t) *t^2

pi = 53.41 * t

t= pi /  53.41

\alpha = 53.41 / ( pi /  53.41)

= 53.41^2 / pi

T = I \alpha = 0.005795 * 908.020 =5.2619 N.m

answer